# Diffusion-based Surrogate Model for Time-varying Underwater Acoustic Channels

### v1

@@ I Introduction
Underwater acoustic communication (UAC) is critical for applications such as ocean exploration, environmental monitoring, defence, and operation of autonomous underwater vehicles [1]. Reliable UAC remains challenging because underwater channels exhibit severe multipath propagation, frequency-selective fading, limited bandwidth, and strong Doppler effects. These characteristics depend on environmental factors such as bathymetry, surface conditions, and sound speed profiles, resulting in dynamic, non-stationary, and time-varying impulse responses (TVIRs) [2]. A TVIR describes the temporal evolution of a channel’s multipath structure, capturing both short-term fading dynamics and long-term statistical variability. Therefore, TVIRs provide a faithful representation of UAC channels. Accurate modeling of TVIRs is useful in the design, evaluation, and deployment of robust underwater communication systems. Furthermore, with the increasing adoption of machine learning (ML) in underwater acoustics, realistic channel models are necessary for generating training datasets for tasks such as localization [3] and adaptive transmission, both of which critically depend on capturing the variability and dynamics of real-world channels.
@@
Traditional approaches to UAC modeling can be broadly categorized into physics methods [4] and replay methods [5]. Physics methods, such as ray-tracing [6] and normal-mode methods [7], provide approximate solutions to the acoustic wave equation under specific assumptions regarding frequency, propagation range, or boundary conditions. Ray-tracing treats acoustic propagation as geometric rays, which is accurate at high frequencies, while normal-mode methods decompose the wavefield into discrete modes, making them effective in low-frequency scenarios. Although these models are physically interpretable and capable of producing high-fidelity outputs when provided with accurate inputs, they depend on detailed environmental parameters (e.g., bathymetry, sound-speed profiles, and seabed properties) and can be computationally intensive. In practice, the difficulty of acquiring precise oceanographic data limits their applicability for channel modeling. Furthermore, these models do not model the time-variability of the channels directly, and require some statistical models of time variability to be adopted to generate TVIRs. By contrast, replay methods reconstruct channel dynamics by reusing measured TVIRs to produce new realizations that preserve key statistical characteristics, including delay spread, Doppler, and path-dependent fading. Common approaches include direct replay [8], which deterministically reproduces subsequent channel responses from measured probes, and stochastic replay [9, 10], which introduces controlled randomness to generate statistically consistent variations beyond the original measurements. Although replay methods achieve high physical fidelity without requiring detailed environmental knowledge, they inherently lack generalization: they cannot extrapolate to unobserved environments, adapt to new deployment scenarios, or generate channel realizations beyond the diversity captured in the available measurements. Consequently, repeated and costly measurements are often required for each new environment.
@@ III-A Overview
Generative models have shown great promise for learning complex, high-dimensional data distributions. Among common paradigms, GANs can produce sharp and realistic samples, making them effective for high-fidelity synthesis [13, 14, 15]. However, GANs are prone to training instability and mode collapse, which can limit sample diversity. VAEs, on the other hand, offer stable training and a probabilistic, interpretable latent space, but standard VAEs often produce slightly blurred outputs and may fail to capture fine-grained temporal or spatial structures [16, 17]. Diffusion models address these limitations by generating data through an iterative denoising process, combining stable training, high-fidelity generation, and the ability to capture complex, multimodal distributions [15, 18]. Although computationally more expensive due to iterative sampling, they are particularly effective for high-dimensional, temporally correlated, and non-stationary data such as TVIRs, overcoming the trade-offs of both GANs and VAEs.
@@ III-B Autoencoder
The autoencoder is designed to compress TVIRs while preserving essential temporal and structural features. TVIRs consist of sequences of CIRs that exhibit strong correlation across consecutive snapshots. Moreover, UAC channels are typically sparse, with only a few significant multipath components present at each time instant. These properties of sparsity and temporal correlation facilitate effective compression. To exploit this, both the encoder and decoder employ Long Short-Term Memory (LSTM) networks, which are well-suited for modeling sequential data with long-term dependencies [20].
@@
Although complex-valued neural networks can naturally model the relationships between amplitude and phase of complex inputs, they are less explored [21], harder to train, and not natively supported in popular frameworks such as PyTorch or TensorFlow. To address this, we convert complex-valued TVIRs $\mathbf{X}$ into a real-valued form suitable for standard LSTM networks. Recall Euler’s formula, which represents a complex number as:
@@ III-C Conditional Latent Diffusion Model
Diffusion models are a state-of-the-art generative framework for synthesizing high-fidelity and diverse samples [15, 18, 22]. These models consist of two complementary processes: a forward process that gradually adds Gaussian noise to a sample until it becomes indistinguishable from isotropic noise, and a reverse process, parameterized by a neural network, which iteratively denoises the corrupted sample to reconstruct the original data [23]. Fig. 3 illustrates these processes applied to a TVIR.
@@
where $\boldsymbol{\epsilon}_{t-1}\sim\mathcal{N}(0,\mathbf{I})$, $t=1,\dots,T$ indexes the diffusion time steps, and $\mathbf{I}$ denotes the identity matrix. The sequence $\{\beta_{t}\}_{t=1}^{T}$ defines the noise schedule [24], which controls the magnitude of Gaussian perturbations added at each step.
@@
The core of the reverse diffusion process is the denoising network, which iteratively removes noise to recover the clean latent representation. In our framework, this network $\kappa(\cdot)$ predicts the noise component at each time step $t$, taking as input the noisy latent $\mathbf{z}_{t}$, a 32-dimensional positional embedding of $t$ [25, 15], and a conditional latent $\mathbf{z}_{c}$ derived from an observed TVIR. This design allows the model to generate realistic, condition-guided channel responses.
@@ IV-A Pre-training on Simulated Dataset
For each simulated environment defined by environmental parameters randomly sampled from the specified ranges, we used the RaySolver from the UnderwaterAcoustics.jl package [26, 27] to generate a nominal CIR. Fig. 5 illustrates the ray trajectories for a representative simulated underwater environment. To introduce temporal variations and obtain a full TVIR from the nominal CIR, we employed an acoustic channel simulator [28], which perturbs the nominal CIR using randomized, time-varying environmental parameters to emulate dynamic underwater conditions. An example nominal CIR generated by RaySolver and the corresponding TVIR produced by the simulator are shown in Fig. 6.
@@ IV-B Autoencoder Performance
We pre-trained the LSTM autoencoder on clean simulated TVIRs to capture essential channel characteristics. The training set comprised 900,000 simulated environments, with 99,500 reserved for validation and 500 for testing. The initial learning rate was set to 0.001 and reduced by a factor of 10 if the validation error did not improve for three consecutive epochs. Early stopping was applied when the learning rate fell below $10^{-6}$. Training was performed using the ADAM optimizer [29] and typically completed within a few hours on an NVIDIA RTX A6000 GPU. Fig. 7 presents the reconstruction results on the clean test set, demonstrating that the autoencoder effectively captures key amplitude variations and generalizes well to unseen environments under temporal variability.
@@ IV-C Diffusion Model Performance
The conditioned and target TVIRs were encoded into latent space using the pre-trained LSTM autoencoder. The diffusion model was then trained to predict latent sequences that, when decoded, reproduced TVIRs consistent with the target sequences. A total of 900,000 conditioned–ground truth latent pairs were used for training, with 99,500 reserved for validation and 500 for testing. Training began with a learning rate of 0.001, which was reduced by a factor of 10 if the validation loss did not improve for five consecutive epochs. Early stopping was applied once no further improvement was observed and the learning rate dropped below $10^{-6}$. The model employed 100 denoising steps with a linear noise schedule ranging from $10^{-4}$ to $10^{-2}$. Although alternative schedules have been explored in prior studies [24, 30], this simple configuration was used for pre-training, and the scheduler was treated as a tunable hyperparameter during fine-tuning, enabling task-specific optimization for improved denoising performance. Pre-training typically completed within one to two hours on an NVIDIA RTX A6000 GPU.
@@ IV-D Experimental Fine-tuning
Measured TVIRs were extracted from the raw recordings using the normalized least-mean-squares (NLMS) adaptive filtering algorithm [31] implemented in AdaptiveEstimators.jl [32]. Fig. 9 shows two example channel responses from different recordings, illustrating delay variations caused by the drifting transmitter. To facilitate model training, amplitude and index normalization was applied to align the recordings. Specifically, each TVIR segment was scaled so that the maximum amplitude of the first CIR equaled one, and its peak was aligned to delay index 20, with subsequent CIRs shifted accordingly. This straightening preserved the spatio-temporal evolution of the channel taps within the TVIR window of interest while providing the model with consistently aligned inputs that were easier to learn from. Fig. 10 illustrates this procedure by stacking straightened TVIRs from consecutive recordings, demonstrating how the normalization removed delay drift while retaining the temporal dynamics of the channel.
@@ V-A Watermark Norway Oslofjord (NOF1) Channel Dataset
This strategy ensures that the diffusion model learns to generate realistic TVIRs tailored to the target environment. We employed a sigmoid-shaped noise scheduler, which allows the model to first focus on fine-grained latent features during the early diffusion steps and progressively handle increasingly randomized latent states in later steps. The sigmoid-shaped noise schedule [34] is defined as:
@@ V-B Keppel Marina Dataset
The discrepancies observed in stochastic replay stem from its underlying assumptions: (i) the UAC channel can be decomposed into a slow-varying trend (large-scale fading) and a fast-varying component (small-scale multipath fading), and (ii) once the slow trend is removed, the fast-varying component is assumed to be stationary with consistent statistics throughout the measurement period. These assumptions hold only if the trend and stationary components can be reliably separated and their statistics remain stable across the observed TVIRs. To achieve this decomposition, stochastic replay employs empirical mode decomposition (EMD) [36] to extract the slow- and fast-fading terms. However, for short-duration TVIRs, as in our Keppel Marina dataset, this process can be inaccurate, introducing artifacts that deviate from the true channel behaviour. This highlights a broader limitation of the stochastic replay method: its effectiveness strongly depends on the validity of the stationarity and decomposition assumptions.

### v2

@@ I Introduction
Underwater acoustic communication (UAC) is critical for applications such as ocean exploration, environmental monitoring, defence, and operation of autonomous underwater vehicles [lurton2002introduction]. Reliable UAC remains challenging because underwater channels exhibit severe multipath propagation, frequency-selective fading, limited bandwidth, and strong Doppler effects. These characteristics depend on environmental factors such as bathymetry, surface conditions, and sound speed profiles, resulting in dynamic, non-stationary, and time-varying impulse responses (TVIRs) [stojanovic2009underwater]. A TVIR describes the temporal evolution of a channel’s multipath structure, capturing both short-term fading dynamics and long-term statistical variability. Therefore, TVIRs provide a faithful representation of UAC channels. Accurate modeling of TVIRs is useful in the design, evaluation, and deployment of robust underwater communication systems. Furthermore, with the increasing adoption of machine learning (ML) in underwater acoustics, realistic channel models are necessary for generating training datasets for tasks such as localization [yuan2025research] and adaptive transmission, both of which critically depend on capturing the variability and dynamics of real-world channels.
@@
Traditional approaches to UAC modeling can be broadly categorized into physics methods [etter2018underwater] and replay methods [otnes2013validation]. Physics methods, such as ray-tracing [gul2017underwater] and normal-mode methods [westwood1996normal], provide approximate solutions to the acoustic wave equation under specific assumptions regarding frequency, propagation range, or boundary conditions. Ray-tracing treats acoustic propagation as geometric rays, which is accurate at high frequencies, while normal-mode methods decompose the wavefield into discrete modes, making them effective in low-frequency scenarios. Although these models are physically interpretable and capable of producing high-fidelity outputs when provided with accurate inputs, they depend on detailed environmental parameters (e.g., bathymetry, sound-speed profiles, and seabed properties) and can be computationally intensive. In practice, the difficulty of acquiring precise oceanographic data limits their applicability for channel modeling. Furthermore, these models do not model the time-variability of the channels directly, and require some statistical models of time variability to be adopted to generate TVIRs. By contrast, replay methods reconstruct channel dynamics by reusing measured TVIRs to produce new realizations that preserve key statistical characteristics, including delay spread, Doppler, and path-dependent fading. Common approaches include direct replay [otnes2013validation], which deterministically reproduces subsequent channel responses from measured probes, and stochastic replay [socheleau2015stochastic, socheleau2015parametric], which introduces controlled randomness to generate statistically consistent variations beyond the original measurements. Although replay methods achieve high physical fidelity without requiring detailed environmental knowledge, they inherently lack generalization: they cannot extrapolate to unobserved environments, adapt to new deployment scenarios, or generate channel realizations beyond the diversity captured in the available measurements. Consequently, repeated and costly measurements are often required for each new environment.
@@ III-A Overview
Generative models have shown great promise for learning complex, high-dimensional data distributions. Among common paradigms, GANs can produce sharp and realistic samples, making them effective for high-fidelity synthesis [goodfellow2020generative, creswell2018generative, wang2017generative]. However, GANs are prone to training instability and mode collapse, which can limit sample diversity. VAEs, on the other hand, offer stable training and a probabilistic, interpretable latent space, but standard VAEs often produce slightly blurred outputs and may fail to capture fine-grained temporal or spatial structures [kingma2019introduction, pinheiro2021variational]. Diffusion models address these limitations by generating data through an iterative denoising process, combining stable training, high-fidelity generation, and the ability to capture complex, multimodal distributions [ho2020denoising, cao2024survey]. Although computationally more expensive due to iterative sampling, they are particularly effective for high-dimensional, temporally correlated, and non-stationary data such as TVIRs, overcoming the trade-offs of both GANs and VAEs.
@@ III-B Autoencoder
The autoencoder is designed to compress TVIRs while preserving essential temporal and structural features. TVIRs consist of sequences of CIRs that exhibit strong correlation across consecutive snapshots. Moreover, UAC channels are typically sparse, with only a few significant multipath components present at each time instant. These properties of sparsity and temporal correlation facilitate effective compression. To exploit this, both the encoder and decoder employ Long Short-Term Memory (LSTM) networks, which are well-suited for modeling sequential data with long-term dependencies [greff2016lstm].
@@
Although complex-valued neural networks can naturally model the relationships between amplitude and phase of complex inputs, they are less explored [lee2022complex], harder to train, and not natively supported in popular frameworks such as PyTorch or TensorFlow. To address this, we convert complex-valued TVIRs $\mathbf{X}$ into a real-valued form suitable for standard LSTM networks. Recall Euler’s formula, which represents a complex number as:
@@ III-C Conditional Latent Diffusion Model
Diffusion models are a state-of-the-art generative framework for synthesizing high-fidelity and diverse samples [ho2020denoising, cao2024survey, yang2023diffusion]. These models consist of two complementary processes: a forward process that gradually adds Gaussian noise to a sample until it becomes indistinguishable from isotropic noise, and a reverse process, parameterized by a neural network, which iteratively denoises the corrupted sample to reconstruct the original data [rombach2022high]. Fig. 3 illustrates these processes applied to a TVIR.
@@
where $\boldsymbol{\epsilon}_{t-1}\sim\mathcal{N}(0,\mathbf{I})$, $t=1,\dots,T$ indexes the diffusion time steps, and $\mathbf{I}$ denotes the identity matrix. The sequence $\{\beta_{t}\}_{t=1}^{T}$ defines the noise schedule [chen2023importance], which controls the magnitude of Gaussian perturbations added at each step.
@@
The core of the reverse diffusion process is the denoising network, which iteratively removes noise to recover the clean latent representation. In our framework, this network $\kappa(\cdot)$ predicts the noise component at each time step $t$, taking as input the noisy latent $\mathbf{z}_{t}$, a 32-dimensional positional embedding of $t$ [vaswani2017attention, ho2020denoising], and a conditional latent $\mathbf{z}_{c}$ derived from an observed TVIR. This design allows the model to generate realistic, condition-guided channel responses.
@@ IV-A Pre-training on Simulated Dataset
For each simulated environment defined by environmental parameters randomly sampled from the specified ranges, we used the RaySolver from the UnderwaterAcoustics.jl package [UnderwaterAcoustics, chitre2023ua] to generate a nominal CIR. Fig. 5 illustrates the ray trajectories for a representative simulated underwater environment. To introduce temporal variations and obtain a full TVIR from the nominal CIR, we employed an acoustic channel simulator [qarabaqi2013acoustic], which perturbs the nominal CIR using randomized, time-varying environmental parameters to emulate dynamic underwater conditions. An example nominal CIR generated by RaySolver and the corresponding TVIR produced by the simulator are shown in Fig. 6.
@@ IV-B Autoencoder Performance
We pre-trained the LSTM autoencoder on clean simulated TVIRs to capture essential channel characteristics. The training set comprised 900,000 simulated environments, with 99,500 reserved for validation and 500 for testing. The initial learning rate was set to 0.001 and reduced by a factor of 10 if the validation error did not improve for three consecutive epochs. Early stopping was applied when the learning rate fell below $10^{-6}$. Training was performed using the ADAM optimizer [kingma2014adam] and typically completed within a few hours on an NVIDIA RTX A6000 GPU. Fig. 7 presents the reconstruction results on the clean test set, demonstrating that the autoencoder effectively captures key amplitude variations and generalizes well to unseen environments under temporal variability.
@@ IV-C Diffusion Model Performance
The conditioned and target TVIRs were encoded into latent space using the pre-trained LSTM autoencoder. The diffusion model was then trained to predict latent sequences that, when decoded, reproduced TVIRs consistent with the target sequences. A total of 900,000 conditioned–ground truth latent pairs were used for training, with 99,500 reserved for validation and 500 for testing. Training began with a learning rate of 0.001, which was reduced by a factor of 10 if the validation loss did not improve for five consecutive epochs. Early stopping was applied once no further improvement was observed and the learning rate dropped below $10^{-6}$. The model employed 100 denoising steps with a linear noise schedule ranging from $10^{-4}$ to $10^{-2}$. Although alternative schedules have been explored in prior studies [chen2023importance, hang2024improved], this simple configuration was used for pre-training, and the scheduler was treated as a tunable hyperparameter during fine-tuning, enabling task-specific optimization for improved denoising performance. Pre-training typically completed within one to two hours on an NVIDIA RTX A6000 GPU.
@@ IV-D Experimental Fine-tuning
Measured TVIRs were extracted from the raw recordings using the normalized least-mean-squares (NLMS) adaptive filtering algorithm [haykin2002adaptive] implemented in AdaptiveEstimators.jl [adaptiveestimation1]. Fig. 9 shows two example channel responses from different recordings, illustrating delay variations caused by the drifting transmitter. To facilitate model training, amplitude and index normalization was applied to align the recordings. Specifically, each TVIR segment was scaled so that the maximum amplitude of the first CIR equaled one, and its peak was aligned to delay index 20, with subsequent CIRs shifted accordingly. This straightening preserved the spatio-temporal evolution of the channel taps within the TVIR window of interest while providing the model with consistently aligned inputs that were easier to learn from. Fig. 10 illustrates this procedure by stacking straightened TVIRs from consecutive recordings, demonstrating how the normalization removed delay drift while retaining the temporal dynamics of the channel.
@@ V-A Watermark Norway Oslofjord (NOF1) Channel Dataset
This strategy ensures that the diffusion model learns to generate realistic TVIRs tailored to the target environment. We employed a sigmoid-shaped noise scheduler, which allows the model to first focus on fine-grained latent features during the early diffusion steps and progressively handle increasingly randomized latent states in later steps. The sigmoid-shaped noise schedule [guo2025comprehensive] is defined as:
@@ V-B Keppel Marina Dataset
The discrepancies observed in stochastic replay stem from its underlying assumptions: (i) the UAC channel can be decomposed into a slow-varying trend (large-scale fading) and a fast-varying component (small-scale multipath fading), and (ii) once the slow trend is removed, the fast-varying component is assumed to be stationary with consistent statistics throughout the measurement period. These assumptions hold only if the trend and stationary components can be reliably separated and their statistics remain stable across the observed TVIRs. To achieve this decomposition, stochastic replay employs empirical mode decomposition (EMD) [huang1998empirical] to extract the slow- and fast-fading terms. However, for short-duration TVIRs, as in our Keppel Marina dataset, this process can be inaccurate, introducing artifacts that deviate from the true channel behaviour. This highlights a broader limitation of the stochastic replay method: its effectiveness strongly depends on the validity of the stationarity and decomposition assumptions.